i need help solving this equation, can you please help me?
solve for n
(4 2/3 + 2 4/7) + 7 3/7 = n + 12

Answer by rfer(16322)

(4 14/21+2 12/21)+7 9/21=n+12
13 35/21=n+12
14 14/21=n+12
14 14/21-12=n
2 2/3=n